# Drawing strategies in Strong Ramsey games for 3-uniform hypergraphs

### v1

@@ 4. $\mathcal{R}(K_{\aleph_{0}}^{(3)},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$
Since $e^{P_{1}}_{T_{2}}(G_{t})=2t+1$, we had $e^{P_{1}}_{T_{1}}(G_{t})=2t$. Moreover, $a_{M}$ has been chosen by $P_{1}$ at latest in her $(2t)^{\textit{th}}$ move, so by choice of $x_{C}$, we have $x_{C}\notin\{a_{1},a_{2},a_{M}\}$. Since $z$ was chosen in move $2t$, we also have $z\notin\{a_{1},a_{2},a_{M}\}$. So $d^{P_{1}}_{T_{4}}(x_{C}),d^{P_{1}}_{T_{4}}(z)\leq t+2$ and since $t\geq 3$, there cannot be a $K_{2,t}^{(3)}$ claimed by $P_{1}$, with center $x_{C}$ or $z$ and hence $\mathit{(ii)}$ and $\mathit{(iii)}$ of the Distraction Lemma (4.1) hold.
If $P_{1}$ had a threat graph $H$ at $T_{4}$, then its center must be contained in at least 3t13t-1 edges. By the same degree argument as above, the only possible choices for its center are $a_{1},a_{2},a_{M}$. The $a_{2}$-board contains a $P_{1}$-star centered at $a_{1}$ and at most one additional $P_{1}$-edge, since $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}\in E_{3}$. Since it could be that $a_{M}=y_{i}$ for some $y_{i}$, the $a_{M}$-board contains at most a $P_{1}$-star and two additional $P_{1}$-edges, so both $a_{2}$ and $a_{M}$ cannot be the centers of $H$. On the $a_{1}$-board, only $a_{2}$ and $a_{M}$ have degree at least 4 at $T_{4}$, so if $t\geq 4$, only $a_{2}$ and $a_{M}$ could be the main vertices of $H$. By definition of $T_{4}$ the last edge $P_{1}$ claimed before $T_{4}$ must be contained in $H$, but cannot be of the form $a_{1}a_{2}v$ or $a_{1}a_{M}v$ for some $v$, because in that case $P_{2}$ reacted with part $\mathit{(i)}$ or $\mathit{(ii)}$ of the algorithm. Therefore, $P_{1}$ cannot have a threat graph at $T_{4}$, if $t\geq 4$.

### v2

@@ 4. $\mathcal{R}(K_{\aleph_{0}}^{(3)},K^{(3)}_{2,t+1}(t-2))$ is a draw for all $t\geq 3$
Since $e^{P_{1}}_{T_{2}}(G_{t})=2t+1$, we had $e^{P_{1}}_{T_{1}}(G_{t})=2t$. Moreover, $a_{M}$ has been chosen by $P_{1}$ at latest in her $(2t)^{\text{th}}$ move, so by choice of $x_{C}$, we have $x_{C}\notin\{a_{1},a_{2},a_{M}\}$. Since $z$ was chosen in move $2t$, we also have $z\notin\{a_{1},a_{2},a_{M}\}$. So $d^{P_{1}}_{T_{4}}(x_{C}),d^{P_{1}}_{T_{4}}(z)\leq t+2$ and since $t\geq 3$, there cannot be a $K_{2,t}^{(3)}$ claimed by $P_{1}$, with center $x_{C}$ or $z$ and hence (ii) and (iii) of the Distraction Lemma (4.1) hold.
If $P_{1}$ had a threat graph $H$ at $T_{4}$, then its center must be contained in at least 3t13t-1 edges. By the same degree argument as above, the only possible choices for its center are $a_{1},a_{2},a_{M}$. The $a_{2}$-board contains a $P_{1}$-star centered at $a_{1}$ and at most one additional $P_{1}$-edge, since $x_{C}zy_{t+1},\dots,x_{C}zy_{2t-1}\in E_{3}$. Since it could be that $a_{M}=y_{i}$ for some $y_{i}$, the $a_{M}$-board contains at most a $P_{1}$-star and two additional $P_{1}$-edges, so both $a_{2}$ and $a_{M}$ cannot be the centers of $H$. On the $a_{1}$-board, only $a_{2}$ and $a_{M}$ have degree at least 4 at $T_{4}$, so if $t\geq 4$, only $a_{2}$ and $a_{M}$ could be the main vertices of $H$. By definition of $T_{4}$ the last edge $P_{1}$ claimed before $T_{4}$ must be contained in $H$, but cannot be of the form $a_{1}a_{2}v$ or $a_{1}a_{M}v$ for some $v$, because in that case $P_{2}$ reacted with part (i) or (ii) of the algorithm. Therefore, $P_{1}$ cannot have a threat graph at $T_{4}$, if $t\geq 4$.